# O90 O1em m #3

### v1

@@ Abstract
Quantum computing is an exciting field with high disruptive potential, but very difficult to access. For this reason, numerous concepts are being developed worldwide on how quantum computing can be taught. This always raises questions about the didactic concept, the content actually taught, and how to measure the success of the teaching concept. In 2022 and 2023, the authors gave a total of nine two-week MOOCs (massive open online courses) with different possible learning paths on the Hasso Plattner Institute’s OpenHPI platform. The platform’s purpose is to make computer science education available to everyone free of charge. The nine quantum courses form a self-contained curriculum. A total of 17157 course attendances have been taken by 7413 natural persons, and the number is still rising. This paper presents the course concept and then evaluates the anonymized data on the background of the participants, their behavior in the courses, and their learning success. In the present paper for the first time such a large dataset of MOOC-based quantum computing education is analyzed. The summarized results are a heterogeneous personal background of the participants biased towards IT professionals, a majority following the didactic recommendations, and a high success rate, which is strongly correlatatd to following the didactic recommendations. The amount of data from such a large group of quantum computing learners offers numerous starting points for further research in the field of quantum computing education.
@@ 1 Introduction
Quantum technologies are considered to be one of the most important technologies of the near future. Recently, the sub-area of quantum computing and quantum communication has increasingly come into focus, see section 2 for a brief glance at the literature. Many companies have launched initial innovation projects in order not to miss out on these technologies. However, a major problem is the lack of skilled workers, particularly in the field of quantum computing. Universities have now set up relevant courses, but the number of graduates will not meet the expected demand for specialists. Politicians have recognized this and have launched various programs. The Federal Ministry of Education and Research in Germany (BMBF) has therefore initiated the ”Quantum Future of Education” funding measure for 2021. The aim was to ”develop innovative, interdisciplinary concepts and programs for education and training in quantum technologies”.
@@
As part of this project, an innovative MOOC (Massive Open Online Courses) curriculum on topics related to quantum computing was developed by the authors of the present paper in cooperation with the Hasso Plattner Institute (HPI). The curriculum consists of nine courses and is explicitly addressed at an audience with little prior knowledge of the subject. The aim was to pick up participants at their level of knowledge and use innovative didactic concepts to convey a sound understanding of various areas of quantum computing. The focus was not on quantum physics or mathematics but on applications of the underlying concepts. As this was a BMBF project and was primarily aimed at a German audience, the MOOC content is in German with an ongoing process of English subtitles being added. The knowledge acquired in the MOOCs is intended to serve as a starting point for further individual specialized professional training.
@@
To make it easier for participants to get started, they are offered various learning paths. This allowes participants to put together an individual learning profile depending on their previous knowledge and interests. Up to January 2024, more than 7413 participants attended the MOOCs. The topic of this paper is to analyze the learning behavior of these participants within the MOOCs. Therefore, the following research questions are addressed:
RQ 1:
Did the MOOCs address a broader audience with different professional and personal characteristics?
RQ 2:
Did the participants follow the given learning paths and other learning recommendations?
RQ 3:
Were the desired competencies acquired by the participants in the courses, and did using the recommended learning items contribute to the learning success?
@@
This paper is organized as follows. Section 2 provides a brief (and necessarily selective) overview of the enormous amount of literature on quantum computing education. Section 3 presents the didactic concept of the curriculum under consideration in this paper, as well as the course’s contents. The content is also categorized with respect to the European Competence Framework for Quantum Technologies [1]. Section 4 evaluates the data collected by HPI during the courses with respect to the three research questions mentioned above. In Section 5, based on the results of section 4, we draw conclusions with respect to the research questions and point out topics for further research.
@@ 2 Literature Review
In contrast to tertiary education, there has been a scant focus on elementary and secondary schools in the realm of educational research on quantum topics. Noteworthy among the few exceptions is the work by Bondani et al. [11]. This study is conducted within the European QTEdu CSA project, a component of the ”Quantum Technologies” flagship, wherein courses on quantum physics and the application of quantum technologies are devised, implemented, and evaluated among approximately 250 students. Freericks et al. [12] already 2017 developed MOOCs for teaching quantum mechanics to over 28.000 non-scientists.
The training imparted in both school and university settings on quantum technologies ultimately aims to address the anticipated global demand for skilled professionals in these domains. This foreseen need is compellingly expounded upon by Venegas-Gomez and Plunkett [13, 14].
@@
Although the aforementioned studies primarily center on the US landscape works such as those by Gerke et al. [17], Greinert et al. [18, 1], contribute to the formulation and presentation of a European competence framework. This framework underpins the development of training programs aimed at nurturing the future quantum workforce. Such initiatives are integral to the European Flagship project QTEdu, which seeks to establish a framework for second-generation quantum technologies.
@@ 3 The OpenHPI Quantum Channel and the curriculum on Quantum Computing
Since 2012, the Hasso Plattner Institute (HPI) operates the online educational platform OpenHPI (https://open.hpi.de). The platform offers free access to MOOCs in the field of information technology. A MOOC at OpenHPI consists usually of around 20 concise videos (approximately 10-15 minutes each), 10 self-tests to assess participant learning progress, and one final examination test. Each course is conducted within a specified timeframe. Participants can repeat self-tests as needed, but the final exam can only be taken once online at the end of the time frame. The time frame is usually two weeks since studies in the literature supported the determination that a two-week module with integrated self-tests and a final exam is optimal [22].
Throughout the course’s timeframe, participants can engage in discussions and pose questions within a forum moderated by the instructor. Answers to questions and comments are usually provided within one day. A course can be completed in two ways. If participants have accessed more than 50% of the items (videos, self-tests, additional material), they receive a certificate of attendance. If participants have more than 50% of the points achieved in the final examination, they receive a qualification certificate.
@@
In the present paper, a distinction between so-called course learners and self-learners is made. Course learners attend the course during the time-frame and are able to interact with the course instructor and fellow students. Self-learners register for the course after the ending of the time-frame. They can only attend the course in a non-interactive way.
In 2022, a dedicated channel for quantum computing was established on OpenHPI (https://open.hpi.de/channels/quantum). The channel features 15 MOOCs addressing various aspects of quantum computing. The kernel is a curriculum of nine interconnected and thematically coordinated modules, developed by the authors of the present paper from June 2022 to July 2023. They consist of three introduction courses, three courses about cryptography, and three courses about algorithms and programming. The last two courses, Algo 2 and Algo 3, were merged into one course for practical reasons. Concurrently, six additional quantum MOOCs were developed as specialized, independent courses. These courses cover diverse topics such as quantum computing for school pupils, a general introduction to Qiskit, or advanced topics like simulating quantum systems (Quantum Computing for Natural Sciences). The present paper discusses and evaluates the curriculum of the nine interconnected quantum MOOCs forming the kernel of the channel.
@@ 3.1 Didactic Concept
The didactic concept aimes to give participants with no prior knowledge of mathematics or physics a sound introduction to quantum information. The broad field is structured by self-contained learning paths so that participants can set their own priorities and not have to deal with topics outside their area of interest. In the end they are able to use the acquired knowledge to face the quantum computing topics occurring in their proper professional practice. This leads to a structure of nine courses. These courses are categorized into three sections: three introductory modules (Intro 1, Intro 2, and Intro 3), three modules focusing on quantum cryptography (Crypto 1, Crypto 2, and Crypto 3), and three modules on quantum algorithms (Algo 1, Algo 2, and Algo 3). The modular structure allows participants to customize their learning paths based on their interests and pre-existing knowledge. The learning paths are illustrated in figure 1.
@@
In the introductory modules, the ”quantum cube” is used throughout to illustrate quantum register states and the effect of quantum gates. This model was developed by Just [23, 24] and since then has been studied and used as an intuitive tool for visualizing quantum entanglement and quantum computation [25, 26, 27]. The quantum cube represents the state of an n-qubit register by providing one dimension in n-dimensional Euclidean space to each qubit. It then utilizes the standard n-dimensional cube with corners $(0,...,0)$ up to $(1,...,1)$. For a given register state, it places the amplitudes of the basis states (in any suitable representation of numbers, i.e. digits, arrows, squares, cubes or circles) at the corresponding corners of the n-dimensional cube. The application of quantum gates to a quantum register can then be visualized as a swapping, rotation, or mixing of the numbers at the corners of the cube. Espescially for dimensions up to 4, quantum entanglement and quantum algorithms can be impressively demonstrated visually, without any prior knowledge of linear algebra. The first two introductory courses use the quantum cube to illustrate simple quantum algorithms like the teleportation algorithm, which as was already published before. In the third introductory course, the cube model is applied for the first time to the advanced issues of quantum oracles and quantum phase kickback. This serves not only to make it easy to understand for the participants but also to demonstrate that the cube model can be applied for visualizing abstract algorithmic concepts [28].
@@
The quantum cryptography series of modules largely dispenses with formal mathematical formulations and uses many examples to illustrate the concepts and mechanisms. For example, simulation examples and graphical diagrams demonstrate entanglement swapping and purification very well.
@@
In the algorithmic course series, the focus is on the practical implementation and less on analytical investigations of the algorithms. Access to the quantum algorithms via programming has the advantage that complex calculations can be omitted and a strong emphasis can be placed on demonstrations.
@@
In addition to the MOOCs, twenty participants were selected by the HPI to participate in a workshop associated with the courses. The workshop was guided by members of the HPI Academy and the authors of the present paper. It used the design thinking method (for further information see e.g. [30]) and aimed to develop tangible applications of quantum technologies. The workshop took the style of a business game. A fictitious company had to be made ”quantum-ready” by the participants. There were no limits to the imagination. Since only a few selected participants could attend the workshop, its evaluation is not the topic of the present paper.
3.2 Module Content
@@
The biggest challenge for the content design of the courses was the different and heterogeneous prior knowledge of the participants. As the courses are aimed at a broad audience, only minimal mathematical and physical knowledge could be required. On the other hand, the aim of the course series was for participants to have a deeper understanding and overview of current topics after successfully completing the curriculum. The bridge between minimal prior knowledge and deeper understanding at the end of the curriculum was built by choosing the contents of the courses as follows.
@@
In the introductory module series, the first module (Intro 1) introduces logical qubits, the elementary gates X, Z, H, and CNOT, and the gate model for the description of algorithms. Since the cube model for illustration [23, 24] is used, this very quickly familiarized the participants with quantum phenomena known from the press such as superposition and entanglement. In the second course (Intro 2), the algorithm for teleportation is presented. Then elementary mathematical description tools ($2^{n}\times 2^{n}$-matrices, the operations of matrix multiplication and tensoring, and complex numbers) are introduced. Formal definitions are deliberately omitted. The focus is on examples and applications to give the participants an intuitive understanding of the concepts. The third module (Intro 3) deals with quantum oracles and the phenomenon of phase kickback, again taking advantage of the quantum cube for visualizing abstract algorithmic concepts. These two concepts form the basis of many classical quantum algorithms. A brief excursion into adiabatic quantum computing concludes the introductory module series.
@@
In the module series on quantum cryptography, the first module (Crypto 1) explains elementary concepts from classical cryptography using examples. The somewhat inaccessible terms computationally secure and perfectly secure are also introduced. It is shown how quantum computers attack common asymmetric encryption methods. The BB84 key exchange protocol [31] is discussed in detail (including simple error correction and privacy amplification)) and its physical implementation and attack possibilities are explained. In the second module (Crypto 2) the central topic is the property of entangled quantum systems (in particular bipartite systems, Schmidt decomposition, partial trace) and the key exchange protocol derived from this. Instead of mathematical calculations, simulations with Qiskit [29] or quantum games (CHSH-Inequality) [32, 33] are used. The ideas on the security proofs of the protocols are also explained. The third module (Crypto 3) discusses the concept of a quantum internet and the associated technological challenges (repeaters and error correction).
[…]

### v2

@@ Abstract
Quantum computing is an exciting field with high disruptive potential, but very difficult to access. For this reason, many approaches to teaching quantum computing are being developed worldwide. This always raises questions about the didactic concept, the content actually taught, and how to measure the success of the teaching concept. In 2022 and 2023, the authors taught a total of nine two-week MOOCs (massive open online courses) with different possible learning paths on the Hasso Plattner Institute’s OpenHPI platform. The purpose of the platform is to make computer science education available to everyone free of charge. The nine quantum courses form a self-contained curriculum. A total of more than 17,000 course attendances have been taken by about 7400 natural persons, and the number is still rising. This paper presents the course concept and evaluates the anonymized data on the background of the participants, their behaviour in the courses, and their learning success. This paper is the first to analyze such a large dataset of MOOC-based quantum computing education. The summarized results are a heterogeneous personal background of the participants biased towards IT professionals, a majority following the didactic recommendations, and a high success rate, which is strongly correlatated with following the didactic recommendations. The amount of data from such a large group of quantum computing learners provides many avenues for further research in the field of quantum computing education. The analyses show that the MOOCs are a low-threshold concept for getting into quantum computing. It was very well received by the participants. The concept can serve as an entry point and guide for the design of quantum computing courses.
@@ 1 Introduction
Quantum technologies are considered to be one of the most important technologies of the near future. Recently, the sub-area of quantum computing and quantum communication has increasingly come into focus, see Sect. 2 for a brief glance at the literature. Many companies have launched initial innovation projects in order not to miss out on these technologies. However, a major problem is the lack of skilled workers, particularly in the field of quantum computing. Universities have now set up relevant courses, but the number of graduates will not meet the expected demand for specialists. Politicians have recognized this and have launched various programs. The Federal Ministry of Education and Research in Germany (BMBF) has therefore initiated the “Quantum Future of Education” funding measure for 2021. The aim was to “develop innovative, interdisciplinary concepts and programs for education and training in quantum technologies”.
@@
As part of this project, an innovative MOOC (Massive Open Online Courses) curriculum on topics related to quantum computing was developed by the authors of the present paper in cooperation with the Hasso Plattner Institute (HPI). The curriculum consists of nine courses and is explicitly addressed to an audience with little prior knowledge of the subject. The aim was to pick up participants at their level of knowledge and use innovative didactic concepts to convey a sound understanding of various areas of quantum computing. The focus was not on quantum physics or mathematics but on applications of the underlying concepts. As this was a BMBF project and was primarily aimed at a German audience, the MOOC content is in German with an ongoing process of English subtitles being added. The knowledge acquired in the MOOCs is intended to serve as a starting point for further individual specialized professional training.
@@
To make it easier for participants to get started, they are offered various learning paths. This allows participants to put together an individual learning profile depending on their previous knowledge and interests. On January 17th, 2024, when the data for this paper was exported, 7413 participants had attended the MOOCs, who had taken an overall of 17,157 courses in the curriculum. The topic of this paper is to analyze the learning behavior of these participants within the MOOCs. Therefore, the following research questions are addressed:
  1. [RQ 2:]
  2. RQ 1:
    What kind of audience attended the MOOCs? What are the specific backgrounds and interests of the participants?
  3. RQ 2:
    What specific topics or subject areas did the participants choose when selecting the courses offered? Which topics are the participants particularly interested in?
  4. RQ 3:
    Which competences could be acquired by the participants through the course concept? To what extent did the course concept support the participants?
@@ 4. RQ 3:
This paper is organized as follows. Section 2 provides a brief (and necessarily selective) overview of the enormous amount of literature on quantum computing education. Section 3 presents the didactic concept of the curriculum under consideration in this paper, as well as the course’s contents. The content is also categorized with respect to the European Competence Framework for Quantum Technologies [1]. Section 4 evaluates the data collected by HPI during the courses with respect to the three research questions mentioned above. In Sect. 5, based on the results of Sect. 4, we draw conclusions concerning the research questions, the lessons learned, and point out topics for further research.
@@ 2 Literature review
In contrast to tertiary education, educational research on quantum topics focussing on elementary and secondary schools just started a few years ago.
The study of Bondani et al. [11] is initiated by the European QTEdu CSA project, a component of the “Quantum Flagship”, wherein courses on quantum physics and the application of quantum technologies are devised, implemented, and evaluated among approximately 250 students. Freericks et al. [12] already 2017 developed MOOCs for teaching quantum mechanics to over 28.000 non-scientists.
The work by Angara et al. [13], Santanassi et al. [14, 15], Pospiech [16], Sutrini et al. [17] investigate the possibilities to include basic quantum algorithms (e.g. teleportation) into physics classes. In addition, Sutirini et al. [17, 18] investigated how training courses for physics teachers can be set up.
The training in both school and university settings on quantum technologies ultimately aims to address the anticipated global demand for skilled professionals in these domains. This foreseen need is compellingly expounded upon by Venegas-Gomez and Plunkett [19, 20].
@@
Although the aforementioned studies primarily center on the US landscape works such as those by Greinert (nee Gerke) et al. [23, 24, 1], contribute to the formulation and presentation of a European competence framework. This framework underpins the development of training programs aimed at nurturing the future quantum workforce. Such initiatives were integral to the European project QTEdu CSA, which began to establish a framework for second-generation quantum technologies - activities that are now continuing as part of QUCATS, the current Quantum Flagship CSA.
@@ 3 The OpenHPI Quantum Channel and the curriculum on Quantum Computing
Since 2012, the Hasso Plattner Institute (HPI) has operated the online educational platform OpenHPI (https://open.hpi.de). The platform offers free access to MOOCs in the field of information technology. A MOOC at OpenHPI usually consists of around 20 concise videos (approximately 10-15 minutes each), 10 self-tests to assess participant learning progress and one final examination test. The self-tests and the final examination tests are carried out online and are multiple-choice tests with more than one choice possible for the questions. The self-tests consist of about 10 questions each, the final exam of 20 questions.
Each course is conducted within a specified timeframe. Participants can repeat self-tests as needed, but the final exam can only be taken once online at the end of the time frame and has to be finished within 120 minutes.
The time frame of a MOOC normally is two weeks since studies in the literature support the determination that a two-week module with integrated self-tests and a final exam is optimal [28].
Throughout the course’s timeframe, participants can engage in discussions and pose questions within a forum moderated by the instructor. Answers to questions and comments are usually provided within one day. A course can be completed in two ways. If participants have accessed more than 50% of the items (videos, self-tests, additional material), they receive a certificate of attendance. If participants have more than 50% of the points achieved in the final examination, they receive a qualification certificate. It should be noted that it is not necessary to access any of the items to obtain a qualification certificate. It is sufficient to pass the final exam. These are the commen rules for all the courses at openHPI.
@@
In the present paper, a distinction between so-called course learners and self-learners is made. Course learners attend the course during the timeframe and are able to interact with the course instructor and fellow students. Self-learners register for the course after the end of the timeframe. They can only attend the course in a non-interactive way.
In 2022, a dedicated channel for quantum computing was established on OpenHPI (https://open.hpi.de/channels/quantum). The channel features 15 MOOCs addressing various aspects of quantum computing. The kernel is an introductory curriculum of nine interconnected and thematically coordinated MOOCs developed by the authors of the present paper from June 2022 to July 2023. Concurrently, six additional quantum MOOCs were developed as specialized, independent courses. These courses cover diverse topics such as quantum computing for school pupils, a general introduction to Qiskit, or advanced topics like simulating quantum systems (Quantum Computing for Natural Sciences). The present paper discusses and evaluates the curriculum of the nine interconnected quantum MOOCs forming the kernel of the channel.
@@ 3.1 Didactic Concept
The didactic concept aims to give participants with no prior knowledge of mathematics or physics a sound introduction to quantum information. The broad field is structured by three pillars “introduction”, “cryptography” and “algorithms”. However, it is not required to work through these pillars one by one. Instead, a series of self-contained learning paths is offered, so that participants can set their own priorities and not have to deal with topics outside their area of interest. In the end, they are able to use the acquired knowledge to face the quantum computing topics occurring in their proper professional practice.
The learning paths are illustrated in Fig. 1. The nine courses are categorized into three sections: three introductory courses (Intro 1, Intro 2, and Intro 3), three courses focusing on quantum cryptography (Crypto 1, Crypto 2, and Crypto 3), and three courses on quantum algorithms (Algo 1, Algo 2, and Algo 3). Intro 1 does not require any prior knowledge. Intro 2 requires Intro 1, but with Intro 1 alone it is already possible to understand Crypto 1. In general, it is possible to understand a course with the knowledge of the courses prior to it in Fig. 1. For instance, Algo 2 can be understood with the knowledge of Intro 1, Intro 2, Intro 3 and Algo 1. Especially, the pillars about crypto and algorithms are distinct one from the other, so participants who want to deepen their knowledge in quantum cryptography need nothing to learn about (advanced) quantum algorithms, and vice versa. A proposed learning path is a subset of the courses following the arrows in Fig. 1 and with each course also containing its predecessors. Information on the structure of the curriculum and the structure of the learning paths is given in the first video of each course, thus providing an orientation on where the course is located in the curriculum, and also information on where to get prior knowledge if necessary for a more advanced course in the curriculum. When useful, also the last video of a course provides this information as a review and outlook. The modular structure allows participants to customize their learning paths based on their interests and pre-existing knowledge. For course-learners, the learning path for all 9 courses was determined by the order of the courses’ timeframes: Intro 1, Crypto 1, Intro 2, Crypto 2, Algo 1, Intro 3, Crypto 3, Algo 2, Algo 3. (Algo 2 and Algo 3 were merged together for practical reasons). They could of course attend subsets of the courses, but in the timeframes’ order. Self-learners can work through the courses in any order of their choice, but it is recommended to complete the previous courses in Fig. 1 before attending a new course.
@@
Special didactic features in the course pillars are as follows:
  • In the introductory courses, the “quantum cube” is used throughout to illustrate quantum register states and the effect of quantum gates. This model was developed by Just [29, 30] and since then has been studied and used as an intuitive tool for visualizing quantum entanglement and quantum computation [31, 32, 33]. It was applied to the more advanced topics of quantum oracles and quantum phase kickback for the first time in the courses under consideration and is presented in greater detail in the appendix of this paper.
@@
  • The quantum cryptography series of courses largely dispenses with formal mathematical formulations and uses many examples to illustrate the concepts and mechanisms. For example, simulation examples and graphical diagrams demonstrate entanglement swapping and purification very well.
@@
  • In the algorithmic course series, the focus is on the practical implementation and less on analytical investigations of the algorithms. Access to the quantum algorithms via programming has the advantage that complex calculations can be omitted, and a strong emphasis can be placed on demonstrations.
@@
In addition to the MOOCs, twenty participants were selected by the HPI to participate in a workshop associated with the courses. The workshop was guided by members of the HPI Academy and the authors of the present paper. It used the design thinking method (for further information see e.g. [35]) and aimed to develop tangible applications of quantum technologies. The workshop took the style of a business game. A fictitious company had to be made “quantum-ready” by the participants. There were no limits to the imagination. Since only a few selected participants could attend the workshop, its evaluation is not the topic of the present paper.
3.2 Course Content
@@
The biggest challenge for the content design of the courses was the different and heterogeneous prior knowledge of the participants. As the courses are aimed at such an audience, only minimal mathematical and physical knowledge could be required. On the other hand, the aim of the course series was for participants to have a deeper understanding and overview of current topics after successfully completing the curriculum. The bridge between minimal prior knowledge and deeper understanding at the end of the curriculum was built by choosing the contents of the courses as follows.
@@
In the introductory course series, the first course (Intro 1) introduces logical qubits, the elementary gates X, Z, H, and CNOT, and the gate model for the description of algorithms. Since the cube model for illustration [29, 30] (see appendix) is used, this very quickly familiarized the participants with quantum phenomena known from the press such as superposition and entanglement. In the second course (Intro 2), the algorithm for teleportation is presented. Then elementary mathematical description tools ($2^{n}\times 2^{n}$-matrices, the operations of matrix multiplication and tensoring, and complex numbers) are introduced. Formal definitions are deliberately omitted. The focus is on examples and applications to give the participants an intuitive understanding of the concepts. The third course (Intro 3) deals with quantum oracles and the phenomenon of phase kickback, again taking advantage of the quantum cube for visualizing abstract algorithmic concepts. These two concepts form the basis of many classical quantum algorithms. A brief excursion into adiabatic quantum computing concludes the introductory course series.
@@
In the course series on quantum cryptography, the first course (Crypto 1) explains elementary concepts from classical cryptography using examples. The somewhat inaccessible terms computationally secure and perfectly secure are also introduced. It is shown how quantum computers attack common asymmetric encryption methods. The BB84 key exchange protocol [36] is discussed in detail (including simple error correction and privacy amplification) and its physical implementation and attack possibilities are explained. In the second course (Crypto 2) the central topic is the property of entangled quantum systems (in particular bipartite systems, Schmidt decomposition, partial trace) and the key exchange protocol derived from this. Instead of mathematical calculations, simulations with Qiskit [34] or quantum games (CHSH-Inequality) [37, 38] are used. The ideas on the security proofs of the protocols are also explained. The third course (Crypto 3) discusses the concept of a quantum internet and the associated technological challenges (repeaters and error correction).
In the first course of the series on quantum algorithms and programming (Algo 1), the implementation of the gate model with Qiskit [34] is presented. The classic quantum algorithms [39] such as teleportation, Deutsch’s algorithm, Deutsch-Jozsa’s algorithm, and Grover’s algorithm are presented using Qiskit. In the second course (Algo 2), somewhat more sophisticated algorithms such as the quantum Fourier transform, the Shor algorithm [40] and the HHL algorithm [41] are explained step by step using example implementations. The third part (Algo 3) deals with current NISQ algorithms [42], which can also run on existing error-prone hardware. In addition to Monte Carlo simulations, VQE and QAOA for solving optimization problems and algorithms are presented, see e.g. [43].